# Implicit Hypergraph Neural Network

### v1

@@ I Introduction
Hypergraphs [1] are a generalization of graphs to higher-order relationships. Graphs, where edges can only connect two nodes, capture pairwise relationships. On the other hand, hyperedges in hypergraphs are free to connect an arbitrary number of nodes, enabling them to define higher-order relations involving multiple nodes. The increased expressiveness of hypergraphs has found applications in various domains, including healthcare (where multiple patients are housed in a single inpatient room) [2], social networks (many users subscribe to a group/subreddit/channel) [3], bio-informatics [4], and cyber-security [5].
The staggering success of Graph Neural Networks (GCNs)  [6], designed for predictive tasks over graphs, has also inspired convolutional neural networks over hypergraphs [7]. Hypergraph neural networks (HNNs) resemble standard GCNs but pass node-embedding messages over hyperedges instead of regular edges. This approach has enabled a number of applications, including disease incidence prediction [2, 8] and echo chamber detection [9], in the domains mentioned above. However, as generalizations of graph neural networks, HNNs inherit some weaknesses unique to graph neural networks.
@@
One such limitation is the fixed radius of information aggregation. GCNs learn representations of nodes by convolving over neighborhoods of a fixed diameter $k$. Typically, $k$ is smaller than the diameter of the network, hence, GCNs cannot capture long-range dependencies between the nodes that go beyond the radius $k$. Stacking multiple layers of GCNs to capture relationships between nodes far away in the network can increase the information aggregation radius; unfortunately, it also leads to performance degradation [10]. This is because as more neighborhood information is aggregated, the node embeddings appear to smooth out, resembling each other more and more and losing discriminative power.
HNNs also convolve over neighborhoods as described by the hyperedges to learn node representations. Hence, the question of whether HNNs can maintain performance while increasing the convolution radius to capture long-range dependency is valid. Here, we first put this question to the test. Through our experiments, we discovered that stacking multiple layers of HNNs leads to performance degradation similar to GCNs (See Figure 1). Intuitively, this pattern is expected as message passing rounds in HNNs typically involve more nodes than standard GCNs. To demonstrate this, we started by running Hypergraph Neural Network [7] for a fixed radius of 2 on the senate-bills dataset [11, 12, 13] and kept track of the performance. We repeated the same experiment while increasing the convolution radius and noticed a degradation pattern (solid blue line) similar to that of a standard GCNs.
Implicit Graph Neural Networks [14], IGNNs, overcome this issue by learning a fixed point representation of nodes by iterating GCN-like models until convergence. Since the number of iterations can be arbitrarily large, these models can capture information from the entire graph. Empirical evidence shows that these approaches overcome the long-range dependency issues while maintaining performance.
@@ II Related Works
Machine Learning on Hypergraphs: ML on hypergraph has garnered significant attention for its ability to model complex higher-order relationships within data [1]. In this paper, we focus on hypergraph node classification, a fundamental task in hypergraph learning. Feng et al. [7] introduced the first message-passing neural network tailored for hypergraph learning, laying the foundation for subsequent advancements. Yadati et al. [15] proposed a novel transformation technique that maps hypergraphs to conventional graphs, facilitating the application of standard graph neural networks. To better capture the importance of high-cardinality hyperedges and high-degree nodes, Dong et al. [16] develop a normalization scheme that enhances representation learning. Bai et al. [17] further refine message passing in hypergraph neural networks by incorporating attention mechanisms. Expanding on the flexibility of hypergraph learning, Chien et al. [18] introduce a highly general framework that integrates the Set Transformer [19] with hypergraph neural networks, enabling more expressive representation learning. This is then followed by recent works be Wang et al. [20] that applies diffusion operators on hypergraphs and Wang et al. [21] that defines parametrized hypergraph-regularized energy functions and demonstrate how their minimizes serve as learnable node embeddings.
This growing body of research underscores the potential of hypergraph-based models in capturing intricate structural dependencies beyond what traditional graph models can achieve. This has led to the deployment of hypergraphs in several applications [22, 2, 23, 24, 25].
Graph Implicit Models: Implicit models define their outputs using fixed-point equations. Bai et al. [26] proposed an equilibrium model for sequence data, where the solution is obtained from a fixed-point equation. The Implicit Graph Neural Network (IGNN) [14] extends this idea to graph data and demonstrates implicit model can capture long-range dependencies on graphs. Park et al. [27] introduce an equilibrium-based GNN with a linear transition map, ensuring contraction to guarantee the existence and uniqueness of a fixed point. Liu et al. [28] developed an infinite-depth GNN that captures long-range dependencies while bypassing iterative solvers by deriving a closed-form solution. El Ghaoui et al. [29] proposed a general implicit deep learning framework, addressing the well-posedness of implicit models. Meanwhile, Chen et al. [30] formulated their model using the diffusion equation as an equilibrium constraint and solve a convex optimization problem to determine the fixed point. Zhong et al. [31] proposed an implicit model for dynamic graph learning and a bilevel training algorithm, which can efficiently capture the long-range dependency on dynamic graphs.
@@ III Method
TABLE I: Summary of symbols and descriptions.
Symbol Descriptions
$\mathcal{H};~\mathcal{V};$$\mathcal{E}$ hypergraph; node set; hyperedge set
$\mathcal{B}$ sampled batch
$v$; $e$ node; hyperedge
$\mathbf{D}_{v};~\mathbf{D}_{e}$ Node degree matrix; hyperedge degree matrix
$\mathbf{X};~\mathcal{Y}$ feature matrix; label set
$d$ number of feature dimension
$f_{\theta};~g_{\phi}$ label classifier; membership classifier
$\mathbf{H};~\mathbf{Z}$ incidence matrix; embeddings
$\mathbf{L};~\mathbf{A}$ laplacian matrix; adjacency matrix
$\ell;~\sigma$ classification loss; activation function
$b_{\Omega}$ affine transformation
$\|\cdot\|_{op}$ operator norm
$\|\cdot\|_{\infty}$ infinity norm
$\otimes$ Kronecker product
@@ III-B Implicit Hypergraph Neural Network
Graph Implicit Models: Implicit models [29, 32, 14] typically learn representations by finding the fixed point solution to a function $f(\mathbf{Z},\mathbf{X})$, where $f(\cdot)$ is typically a universal function approximator, such as a neural network. Here $\mathbf{X}$ represents the input features, and $\mathbf{Z}$ denotes the learned embeddings. The fixed-point embeddings are obtained iteratively as follows:
@@
Such a function $f(\cdot)$ is said to be well-posed if there is a fixed-point solution to it (i.e., it produces converged representations that do not change on further iteration). To ensure well-posedness, $f(\cdot)$ is typically required to be a contractive mapping [14].
@@
Note that all implicit graph neural networks satisfy the well-posedness property [14].
Hypergraph Neural Network: Any implicit hypergraph neural network we design must utilize a graph neural network specifically designed for hypergraphs. Hence, a straightforward approach for our aspirations is to start with an existing hypergraph neural network and define an implicit function over it. Let us examine one of the state-of-the-art hypergraph neural networks (HGNN) [7], as illustrated in the equation below.
@@
Here, $\mathbf{Z}_{t}$ is the embedding at the $t^{th}$ layer and $\mathbf{W}$ is a learnable parameter. The rest of the notations are as defined earlier. Now, for further discussion let $\mathbf{L}:=\mathbf{D}_{v}^{-\frac{1}{2}}\mathbf{H}\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$. The equation above can now be simplified as $\mathbf{Z}_{t+1}=\sigma(\mathbf{L}\mathbf{Z}_{t}\mathbf{W})$. This formulation treats $\mathbf{L}$, which is the approximated hypergraph Laplacian matrix [7], analogously to the adjacency matrix in conventional graphs. Since the function $\sigma$ is a non-expansive activation function (e.g., ReLU, sigmoid), the formulation leads to a meaningful implicit graph neural network over hypergraphs. However, a key drawback of this formulation is that it solely focuses on learning node embeddings. It disregards the key distinction between hypergraphs and conventional graphs, the varying cardinality of hyperedges. Consequently, developing an implicit model based on Equation 2 can only yield suboptimal results (as validated by our experiments).
@@
For the well-posedness of our model, we provide the following lemma, which has similar requirements as other implicit models [14, 31]
@@
We first rewrite the equation (4) as $Z=\sigma(LZW+B)$. Showing $W$ is well-posed for equation (4) is equivalent to show the following vectorized equation admits a unique solution: $\text{vec}(Z)=\sigma\left((W^{T}\otimes L)\text{vec}(Z)+\text{vec}(B)\right)$. By Lemma B.1 from [14], we know that if $\|W^{T}\otimes L\|_{op}<1$ then the above equation has a unique solution, which can be attained by fixed-point iteration. Moreover, this constraint can be further converted to $\|W\|_{\infty}\|L\|_{op}\leq\|W\|_{op}\|L\|_{op}=\|W^{T}\otimes L\|_{op}<1$. ∎
Membership Regularization: To further capture the relationship between nodes and hyperedges, we introduce a novel membership regularization technique. Inspired by the labeling trick [33], which enhances representation learning by incorporating node membership information, our approach encourages the model to better leverage node-hyperedge interactions. Given fixed-point node embeddings $\mathbf{Z}_{v}$ and hyperedge embeddings $\mathbf{Z}_{e}$, we impose membership regularization as follows:
@@
$\min_{\theta,\mathbf{W},\Omega,\phi}$ $)\displaystyle\mathcal{L}(\theta,\mathbf{W},\Omega,\phi):=\frac{1}{|\mathcal{V}_{\text{train}}|}\sum_{i\in\mathcal{V}_{\text{train}}}\ell_{1}(f_{\theta}(\hat{\mathbf{Z}}),y_{i})+\gamma\ell_{2}(g_{\phi}(\hat{\mathbf{Z}}$)) (5)
s.t. $\hat{\mathbf{Z}}=\sigma\left(\bar{\mathbf{A}}\hat{\mathbf{Z}}\mathbf{W}+b_{\Omega}(\hat{\mathbf{X}})\right),~\|\mathbf{W}\|_{\infty}\leq\frac{\kappa}{\|\bar{\mathbf{A}}\|_{\text{op}}}.$
@@ III-C Training
To optimize for the objective, we adopt a projected gradient descent through implicit differentiation, which is a standard training algorithm for implicit models [14], enabling direct gradient computation through the fixed point. After updating the parameters, $\mathbf{W}$ will be projected back to the feasible set, for which efficient methods exist [34]. This projection is also used in other implicit models, such as [14].
@@ IV-A Datasets
TABLE II: Summary Statistics of the Datasets used
@@
TABLE III: Results for transductive node classification experiments. We report the mean accuracy with standard deviation taken across 10 independent runs. As observed, our proposed approach IHNN performs the best in all datasets.
@@
  • DBLP is a co-authorship dataset [15], where the nodes are authors and the hyperedges represent co-authorship relations. Node labels correspond to any of the 6 conference categories, namely ‘algorithms’, ‘database’, ‘programming’, ‘datamining’, ‘intelligence’, and ‘vision’. For all the experiments, we use the train: test split given in [15]. Note that each split has 70 to 30 train to test ratio.
  • Cora: We use the Cora co-authorship dataset and the train: test splits created by [15]. Here, the predictive task is again a multi-class node classification like DBLP.
  • Walmart-trips: In this dataset [35], the hyperedges are sets of co-purchased products at Walmart, while the nodes denote the products themselves. The node labels denote the 11 broad categories the products, making the predictive task on this dataset also a multi-class node classification task.
  • High-school: This dataset, obtained from [36] is constructed from interactions recorded by wearable sensors by students at a high school. The sensors record interactions at a resolution of 20 seconds. Each hyperedge corresponds to a group of students that were all in proximity of one another at a given time, based on data from sensors worn by students. Each node is labeled with the grade the student belongs to.
  • House-bills: This dataset, obtained from [35, 11, 37] is constructed from the co-sponsorship of bills in the US House of Representatives. The nodes denote US Congress-persons, and the hyperedges denote the co-sponsors of bills put forth in the House of Representatives. Node label denotes the representative’s political party affiliation.
The summary statistics of the datasets are given in Table II.
@@ IV-B Baselines and Evaluation Metric
  • Confidence Interval Method (CI): Here the exact ranges of unlabeled vertices of a hypergraph are estimated through a confidence interval approach to determine the optimal solutions and a simpler subgradient method for solving the convex program [38]. Note that this method only leverages the structure of the hypergraph to obtain the node embeddings.
@@
  • Multilayer Perceptron with Hypergraph Laplacian Regularization (MLP+HLR): This baseline method [39] indirectly imposes the hypergraph structure by regularizing the embeddings learned by the MLP with a regularizer that depends on the hypergraph laplacian.
  • Hypergraph Neural Network (HGNN): This method [7] obtains the node embeddings by defining a heuristic hypergraph convolution module parameterized by the hypergraph Laplacian.
  • Hypergraph Convolution and Hypergraph Attention (HCHA): Building on the message propagation through hypergraph convolution as in HGNN, this method [17] uses different degree normalizations and attention weights, with the attention weights depending on node features and the hyperedge features. However, the absence of explicit hyperedge features in the datasets prevents the application of attention to the hyperedges.
  • Hypergraph Networks with Hyperedge Neurons (HNHN): This is again a variant [16] of the hypergraph convolution combines a hypergraph convolution network with nonlinear activation functions and a normalization scheme to adjust the importance of high-cardinality hyperedges and high-degree vertices.
  • Graph Convolutional Networks on Hypergraphs (HyperGCN): This method [15] proposes a unique transformation method of a hypergraph to a graph and approximates each hyperedge of the hypergraph by a set of pairwise edges connecting the vertices of the hyperedge and treats the learning problem as a graph learning problem on the approximation.
  • AllSet Transformer (Allset): This baseline [18] uses an attention mechanism during the pooling of information during the hypergraph convolution step. However, this comes with an additional cost of a larger number of trainable parameters, which may not be optimized with less training data.
  • Equivariant Hypergraph Diffusion Neural Network(ED-HNN): This baseline [40] uses diffusion mechanisms to design a new hypergraph message passing structure.
  • Purposeful Hyperedges in Optimization Motivated Neural Networks(PhenomNN): This baseline [21] defines parametrized hypergraph-regularized energy functions and demonstrates how their minimizers serve as learnable node embeddings.
Evaluation Metric: Like the prior works [18, 16], we use accuracy as the evaluation metric for all our experiments. While some related work [15] report the mean test error, note that it is simply (100 - average accuracy).

### v2

@@ I Introduction
Hypergraphs [antelmi2023survey] are a generalization of graphs to higher-order relationships. Graphs, where edges can only connect two nodes, capture pairwise relationships. On the other hand, hyperedges in hypergraphs are free to connect an arbitrary number of nodes, enabling them to define higher-order relations involving multiple nodes. The increased expressiveness of hypergraphs has found applications in various domains, including healthcare (where multiple patients are housed in a single inpatient room) [choudhuri2025domain], social networks (many users subscribe to a group/subreddit/channel) [li2013link], bio-informatics [tian2009hypergraph], and cyber-security [lin2024hypergraph].
The staggering success of Graph Neural Networks (GCNs)  [kipf2016semi], designed for predictive tasks over graphs, has also inspired convolutional neural networks over hypergraphs [feng2019hypergraph]. Hypergraph neural networks (HNNs) resemble standard GCNs but pass node-embedding messages over hyperedges instead of regular edges. This approach has enabled a number of applications, including disease incidence prediction [choudhuri2025domain, anand2024h2abm] and echo chamber detection [hickok2022bounded], in the domains mentioned above. However, as generalizations of graph neural networks, HNNs inherit some weaknesses unique to graph neural networks.
@@
One such limitation is the fixed radius of information aggregation. GCNs learn representations of nodes by convolving over neighborhoods of a fixed diameter $k$. Typically, $k$ is smaller than the diameter of the network, hence, GCNs cannot capture long-range dependencies between the nodes that go beyond the radius $k$. Stacking multiple layers of GCNs to capture relationships between nodes far away in the network can increase the information aggregation radius; unfortunately, it also leads to performance degradation [chen2020measuring]. This is because as more neighborhood information is aggregated, the node embeddings appear to smooth out, resembling each other more and more and losing discriminative power.
HNNs also convolve over neighborhoods as described by the hyperedges to learn node representations. Hence, the question of whether HNNs can maintain performance while increasing the convolution radius to capture long-range dependency is valid. Here, we first put this question to the test. Through our experiments, we discovered that stacking multiple layers of HNNs leads to performance degradation similar to GCNs (See Figure 1). Intuitively, this pattern is expected as message passing rounds in HNNs typically involve more nodes than standard GCNs. To demonstrate this, we started by running Hypergraph Neural Network [feng2019hypergraph] for a fixed radius of 2 on the senate-bills dataset [chodrow2021generative, fowler2006legislative] and kept track of the performance. We repeated the same experiment while increasing the convolution radius and noticed a degradation pattern (solid blue line) similar to that of a standard GCNs.
Implicit Graph Neural Networks [gu2020implicit], IGNNs, overcome this issue by learning a fixed point representation of nodes by iterating GCN-like models until convergence. Since the number of iterations can be arbitrarily large, these models can capture information from the entire graph. Empirical evidence shows that these approaches overcome the long-range dependency issues while maintaining performance.
@@ II Related Works
Machine Learning on Hypergraphs: ML on hypergraph has garnered significant attention for its ability to model complex higher-order relationships within data [antelmi2023survey]. In this paper, we focus on hypergraph node classification, a fundamental task in hypergraph learning. Feng et al. [feng2019hypergraph] introduced the first message-passing neural network tailored for hypergraph learning, laying the foundation for subsequent advancements. Yadati et al. [yadati2019hypergcn] proposed a novel transformation technique that maps hypergraphs to conventional graphs, facilitating the application of standard graph neural networks. To better capture the importance of high-cardinality hyperedges and high-degree nodes, Dong et al. [dong2020hnhn] develop a normalization scheme that enhances representation learning. Bai et al. [bai2021hypergraph] further refine message passing in hypergraph neural networks by incorporating attention mechanisms. Expanding on the flexibility of hypergraph learning, Chien et al. [chienyou] introduce a highly general framework that integrates the Set Transformer [lee2019set] with hypergraph neural networks, enabling more expressive representation learning. This is then followed by recent works be Wang et al. [wang2022equivariant] that applies diffusion operators on hypergraphs and Wang et al. [wang2023hypergraph] that defines parametrized hypergraph-regularized energy functions and demonstrate how their minimizes serve as learnable node embeddings.
This growing body of research underscores the potential of hypergraph-based models in capturing intricate structural dependencies beyond what traditional graph models can achieve. This has led to the deployment of hypergraphs in several applications [xu2023hypergraph, choudhuri2025domain, la2022music, wang2021session].
Graph Implicit Models: Implicit models define their outputs using fixed-point equations. Bai et al. [bai2019deep] proposed an equilibrium model for sequence data, where the solution is obtained from a fixed-point equation. The Implicit Graph Neural Network (IGNN) [gu2020implicit] extends this idea to graph data and demonstrates implicit model can capture long-range dependencies on graphs. Park et al. [park2021convergent] introduce an equilibrium-based GNN with a linear transition map, ensuring contraction to guarantee the existence and uniqueness of a fixed point. Liu et al. [liu2021eignn] developed an infinite-depth GNN that captures long-range dependencies while bypassing iterative solvers by deriving a closed-form solution. El Ghaoui et al. [el2021implicit] proposed a general implicit deep learning framework, addressing the well-posedness of implicit models. Meanwhile, Chen et al. [chen2022optimization] formulated their model using the diffusion equation as an equilibrium constraint and solved a convex optimization problem to determine the fixed point. Zhong et al. [zhong2024efficient] proposed an implicit model for dynamic graph learning and a bilevel training algorithm, which can efficiently capture the long-range dependency on dynamic graphs and subgraphs [zhongimplicit].
@@ III-B Implicit Hypergraph Neural Network
Graph Implicit Models: Implicit models [el2021implicit, geng2021training, gu2020implicit] typically learn representations by finding the fixed point solution to a function $f(\mathbf{Z},\mathbf{X})$, where $f(\cdot)$ is typically a universal function approximator, such as a neural network. Here $\mathbf{X}$ represents the input features, and $\mathbf{Z}$ denotes the learned embeddings. The fixed-point embeddings are obtained iteratively as follows:
@@
Such a function $f(\cdot)$ is said to be well-posed if there is a fixed-point solution to it (i.e., it produces converged representations that do not change on further iteration). To ensure well-posedness, $f(\cdot)$ is typically required to be a contractive mapping [gu2020implicit].
@@
Note that all implicit graph neural networks satisfy the well-posedness property [gu2020implicit].
Hypergraph Neural Network: Any implicit hypergraph neural network we design must utilize a graph neural network specifically designed for hypergraphs. Hence, a straightforward approach for our aspirations is to start with an existing hypergraph neural network and define an implicit function over it. Let us examine one of the state-of-the-art hypergraph neural networks (HGNN) [feng2019hypergraph], as illustrated in the equation below.
@@
Here, $\mathbf{Z}_{t}$ is the embedding at the $t^{th}$ layer and $\mathbf{W}$ is a learnable parameter. The rest of the notations are as defined earlier. Now, for further discussion let $\mathbf{L}:=\mathbf{D}_{v}^{-\frac{1}{2}}\mathbf{H}\mathbf{D}_{e}^{-1}\mathbf{H}^{\top}\mathbf{D}_{v}^{-\frac{1}{2}}$. The equation above can now be simplified as $\mathbf{Z}_{t+1}=\sigma(\mathbf{L}\mathbf{Z}_{t}\mathbf{W})$. This formulation treats $\mathbf{L}$, which is the approximated hypergraph Laplacian matrix [feng2019hypergraph], analogously to the adjacency matrix in conventional graphs. Since the function $\sigma$ is a non-expansive activation function (e.g., ReLU, sigmoid), the formulation leads to a meaningful implicit graph neural network over hypergraphs. However, a key drawback of this formulation is that it solely focuses on learning node embeddings. It disregards the key distinction between hypergraphs and conventional graphs, the varying cardinality of hyperedges. Consequently, developing an implicit model based on Equation 2 can only yield suboptimal results (as validated by our experiments).
@@
For the well-posedness of our model, we provide the following lemma, which has similar requirements as other implicit models [gu2020implicit, zhong2024efficient]
@@
We first rewrite the equation (4) as $Z=\sigma(\bar{\mathbf{A}}\mathbf{ZW}+\mathbf{B})$. Showing $W$ is well-posed for equation (4) is equivalent to show the following vectorized equation admits a unique solution: $\text{vec}({\mathbf{Z}})=\sigma\left(({\mathbf{W}}^{T}\otimes\bar{\mathbf{A}})\text{vec}({\mathbf{Z}})+\text{vec}({\mathbf{B}})\right)$. By Lemma B.1 from [gu2020implicit], we know that if $\|\mathbf{W}^{T}\otimes\bar{\mathbf{A}}\|_{op}<1$ then the above equation has a unique solution, which can be attained by fixed-point iteration. Moreover, this constraint can be further converted to $\|{\mathbf{W}}\|_{\infty}\|\bar{\mathbf{A}}\|_{op}\leq\|{\mathbf{W}}\|_{op}\|\bar{\mathbf{A}}\|_{op}=\|{\mathbf{W}}^{T}\otimes\bar{\mathbf{A}}\|_{op}<1$. ∎
Membership Regularization: To further capture the relationship between nodes and hyperedges, we introduce a novel membership regularization technique. Inspired by the labeling trick [zhang2021labeling], which enhances representation learning by incorporating node membership information, our approach encourages the model to better leverage node-hyperedge interactions. Given fixed-point node embeddings $\mathbf{Z}_{v}$ and hyperedge embeddings $\mathbf{Z}_{e}$, we impose membership regularization as follows:
@@
$\min_{\theta,\mathbf{W},\Omega,\phi}\;$ $\mathcal{L}(\theta,\mathbf{W},\Omega,\phi):=\frac{1}{|\mathcal{V}_{\text{train}}|}\sum_{i\in\mathcal{V}_{\text{train}}}\ell_{1}\!\left(f_{\theta}(\hat{\mathbf{Z}}),y_{i}\right)$ (5)
[…]
s.t. $\hat{\mathbf{Z}}=\sigma\!\left(\bar{\mathbf{A}}\hat{\mathbf{Z}}\mathbf{W}+b_{\Omega}(\hat{\mathbf{X}})\right),~\|\mathbf{W}\|_{\infty}\leq\frac{\kappa}{\|\bar{\mathbf{A}}\|_{\text{op}}}.$ (6)
@@ III-C Training
To optimize for the objective, we adopt a projected gradient descent through implicit differentiation, which is a standard training algorithm for implicit models [gu2020implicit], enabling direct gradient computation through the fixed point. After updating the parameters, $\mathbf{W}$ will be projected back to the feasible set, for which efficient methods exist [duchi2008efficient]. This projection is also used in other implicit models, such as [gu2020implicit].
@@ IV-A Datasets
TABLE I: Summary Statistics of the Datasets used
@@
TABLE II: Results for transductive node classification experiments. We report the mean accuracy with standard deviation taken across 10 independent runs. As observed, our proposed approach IHNN performs the best in all datasets.
@@
  • DBLP is a co-authorship dataset [yadati2019hypergcn], where the nodes are authors and the hyperedges represent co-authorship relations. Node labels correspond to any of the 6 conference categories, namely ‘algorithms’, ‘database’, ‘programming’, ‘datamining’, ‘intelligence’, and ‘vision’. For all the experiments, we use the train: test split given in [yadati2019hypergcn]. Note that each split has 70 to 30 train to test ratio.
  • Cora: We use the Cora co-authorship dataset and the train: test splits created by [yadati2019hypergcn]. Here, the predictive task is again a multi-class node classification like DBLP.
  • Walmart-trips: In this dataset [amburg2020clustering], the hyperedges are sets of co-purchased products at Walmart, while the nodes denote the products themselves. The node labels denote the 11 broad categories the products, making the predictive task on this dataset also a multi-class node classification task.
  • High-school: This dataset, obtained from [benson2018simplicial] is constructed from interactions recorded by wearable sensors by students at a high school [amiri2018netgist]. The sensors record interactions at a resolution of 20 seconds. Each hyperedge corresponds to a group of students that were all in proximity of one another at a given time, based on data from sensors worn by students. Each node is labeled with the grade the student belongs to.
  • House-bills: This dataset, obtained from [amburg2020clustering, chodrow2021generative, veldt2020minimizing] is constructed from the co-sponsorship of bills in the US House of Representatives. The nodes denote US Congress-persons, and the hyperedges denote the co-sponsors of bills put forth in the House of Representatives. Node label denotes the representative’s political party affiliation.
The summary statistics of the datasets are given in Table I.
@@ IV-B Baselines and Evaluation Metric
  • Confidence Interval Method (CI): Here the exact ranges of unlabeled vertices of a hypergraph are estimated through a confidence interval approach to determine the optimal solutions and a simpler subgradient method for solving the convex program [zhang2017re]. Note that this method only leverages the structure of the hypergraph to obtain the node embeddings.
@@
  • Multilayer Perceptron with Hypergraph Laplacian Regularization (MLP+HLR): This baseline method [zhou2006learning] indirectly imposes the hypergraph structure by regularizing the embeddings learned by the MLP with a regularizer that depends on the hypergraph laplacian.
  • Hypergraph Neural Network (HGNN): This method [feng2019hypergraph] obtains the node embeddings by defining a heuristic hypergraph convolution module parameterized by the hypergraph Laplacian.
  • Hypergraph Convolution and Hypergraph Attention (HCHA): Building on the message propagation through hypergraph convolution as in HGNN, this method [bai2021hypergraph] uses different degree normalizations and attention weights, with the attention weights depending on node features and the hyperedge features. However, the absence of explicit hyperedge features in the datasets prevents the application of attention to the hyperedges.
  • Hypergraph Networks with Hyperedge Neurons (HNHN): This is again a variant [dong2020hnhn] of the hypergraph convolution combines a hypergraph convolution network with nonlinear activation functions and a normalization scheme to adjust the importance of high-cardinality hyperedges and high-degree vertices.
  • Graph Convolutional Networks on Hypergraphs (HyperGCN): This method [yadati2019hypergcn] proposes a unique transformation method of a hypergraph to a graph and approximates each hyperedge of the hypergraph by a set of pairwise edges connecting the vertices of the hyperedge and treats the learning problem as a graph learning problem on the approximation.
  • AllSet Transformer (Allset): This baseline [chienyou] uses an attention mechanism during the pooling of information during the hypergraph convolution step. However, this comes with an additional cost of a larger number of trainable parameters, which may not be optimized with less training data.
  • Equivariant Hypergraph Diffusion Neural Network(ED-HNN): This baseline [wangequivariant] uses diffusion mechanisms to design a new hypergraph message passing structure.
  • Purposeful Hyperedges in Optimization Motivated Neural Networks(PhenomNN): This baseline [wang2023hypergraph] defines parametrized hypergraph-regularized energy functions and demonstrates how their minimizers serve as learnable node embeddings.
Evaluation Metric: Like the prior works [chienyou, dong2020hnhn], we use accuracy as the evaluation metric for all our experiments. While some related work [yadati2019hypergcn] report the mean test error, note that it is simply (100 - average accuracy).